### v1

# Search for an eV-scale sterile neutrino with the first six detection units of KM3NeT/ORCA

Abstract
The existence of an eV-scale sterile neutrino has been proposed to explain several anomalous experimental results obtained over the course of the past 25 years. The first search for such a sterile neutrino conducted with data from KM3NeT/ORCA — a water Cherenkov neutrino telescope under construction at the bottom of the Mediterranean Sea — is reported in this paper. GeV-scale atmospheric neutrino oscillations are measured by reconstructing the energy and arrival direction of up-going neutrinos that have traversed the Earth. This study is based on a data sample containing 5828 neutrino candidates collected with 6 detection units ($5\%$ of the complete detector), corresponding to an exposure of 433 kton-years. From the expected effect of an eV-scale sterile neutrino on the first $\nu_{\mu}\rightarrow\nu_{\tau}$ standard oscillation maximum, simultaneous constraints are put on the magnitude of the $U_{\mu 4}$ and $U_{\tau 4}$ mixing elements under the assumption $\Delta m^{2}_{41}=1$ eV2. The results are compatible with the absence of mixing between active neutrinos and a sterile state, with $|U_{\mu 4}|^{2}<0.138$ and $|U_{\tau 4}|^{2}<0.076$ at a $90\%$ confidence level. Such constraints are compatible with the results reported by other long-baseline experiments, and indicate that with KM3NeT/ORCA it is possible to bring crucial contributions to sterile neutrino searches in the coming years.
1 Introduction
Since their first observation in 1998 super-kamiokandecollaborationEvidenceOscillationAtmospheric1998 ; ambrosioMeasurementAtmosphericNeutrinoinduced1998 , neutrino oscillations have been modeled through the three-flavour Pontecorvo-Maki-Nakagawa-Sakata (PMNS) paradigm dentonSnowmassNeutrinoFrontier2022 . The model describes the mixing between the neutrino flavour eigenstates ($\nu_{e}$, $\nu_{\mu}$, $\nu_{\tau}$) and the neutrino mass eigenstates ($\nu_{1}$, $\nu_{2}$, $\nu_{3}$) of masses ($m_{1}$, $m_{2}$, $m_{3}$). In this formalism, the PMNS matrix is a 33-by-33 rotation matrix parametrised by the three mixing angles $\theta_{12}$, $\theta_{13}$ and $\theta_{23}$ and the Dirac CP-violating phase $\delta$. The mixing angles determine the extent of mixing between the neutrino states, while the CP-violating phase breaks the symmetry of oscillation probabilities for neutrinos and antineutrinos. The frequencies of the observed oscillations are determined by the values of two independent squared-mass differences, for example $\Delta m^{2}_{31}$ and $\Delta m^{2}_{21}$. Over the past 25 years, experiments using various neutrino sources have significantly improved the precision in measuring these six oscillation parameters, with uncertainties currently down to the percent level estebanFateHintsUpdated2020 . However, the $\delta$ phase is still largely unconstrained by experiments. The $\theta_{23}$ octant (whether it is lower than, equal to or higher than $\pi/4$) is also undetermined. Finally, the sign of $\Delta m^{2}_{31}$, determining the neutrino mass ordering (NMO) which is either normal ($m_{1}<m_{2}<m_{3}$, NO) or inverted ($m_{3}<m_{1}<m_{2}$, IO), is also unknown. Current and upcoming large-scale neutrino oscillation experiments will pursue these questions over the next decade to deepen our understanding of neutrino properties.
Despite the success of the three-flavour PMNS paradigm at describing the neutrino oscillation data, several experimental results obtained at a relatively short baseline (SBL) remain largely unexplained aceroWhitePaperLight2023 . All these anomalies, despite originating from experiments with different neutrino sources, detection technologies, oscillation channels, baselines and neutrino energies, can be individually interpreted as oscillations driven by a squared-mass difference $\Delta m^{2}\sim\mathcal{O}(1)$ eV2. This value of $\Delta m^{2}$ is several orders of magnitude larger than the measured values of $\Delta m^{2}_{31}$ and $\Delta m^{2}_{21}$. Thus, to explain the SBL anomalies in terms of oscillations, a simple extension of the standard model is to assume the existence of a fourth, eV-scale, neutrino state. However, cosmological constraints exist on the effective number of relativistic neutrino species lesgourguesMassiveNeutrinosCosmology2006 ; navasReviewParticlePhysics2024 . In addition, the LEP experiments obtained compelling evidence that only three flavours of light neutrinos that couple to the weak interaction — known as active neutrinos — exist PrecisionElectroweakMeasurements2006 . [^1] In that context, light means with a mass lower than half of the mass of the Z boson. Thus, an additional eV-scale neutrino would have to be insensitive to the weak interaction, or sterile. In the following, this model will be referred to as the 3++1 model.
The presence of an eV-scale sterile neutrino would also alter the oscillation probabilities of neutrinos detected in experiments with longer baselines such as atmospheric neutrino telescopes. The IceCube collaboration recently reported updated results of searches for an eV-scale sterile neutrino using their complete detector icecubecollaborationSearchEVscaleSterile2024a ; abbasiExplorationMassSplitting2024a as well as using the denser DeepCore array icecubecollaborationSearchLightSterile2024 . In the present work, the first search for an eV-scale sterile neutrino performed with data from the KM3NeT/ORCA detector is presented. After a review of the effect of an eV-scale sterile neutrino on the oscillation channels of interest in section 2, the KM3NeT/ORCA water Cherenkov neutrino telescope is presented in section 3. The dataset used in this study was collected with an early partial configuration of the detector with six detection units, called ORCA6 in the following, operating from January 2020 to November 2021. The data sample as well as the event selection strategy is presented in section 4. The model and statistical methods used to perform this oscillation analysis are presented in section 5. Finally, the results are summarised and discussed in sections 6 and 7 respectively.
2 Effect of an eV-scale sterile neutrino on the oscillation probabilities
Oscillations in the presence of a single sterile neutrino can be modeled by extending the standard three-flavour formalism to include a fourth mass eigenstate $m_{4}$. This extension introduces one additional squared-mass difference, usually denoted by $\Delta m_{41}^{2}$. The PMNS matrix becomes a $4\times 4$ unitary matrix $U_{\alpha i}$ with $\alpha=e,\mu,\tau,s$ and $i=1,2,3,4$, which can be parametrised such that:
$U=R_{34}\tilde{R}_{24}\tilde{R}_{14}R_{23}\tilde{R}_{13}R_{12}\,$, (1)
where $R_{jk}$ is a rotation matrix in the $j$-$k$ plane and, similarly, $\tilde{R}_{jk}$ is a unitary rotation matrix with an added complex phase. In addition to the three standard mixing angles and the Dirac CP violation phase, three active-sterile mixing angles $\theta_{i4}$, $i=1,2,3$, and two additional Dirac CP violation phases $\delta_{i4}$, $i=1,2$ are usually introduced to parametrise the extended matrix. The active-sterile mixing angles are related to the active-sterile mixing elements by:
$U_{e4}=\sin\theta_{14}e^{-i\delta_{14}}$ (2)
$U_{\mu 4}=\cos\theta_{14}\sin\theta_{24}e^{-i\delta_{24}}$ (3)
$U_{\tau 4}=\cos\theta_{14}\cos\theta_{24}\sin\theta_{34}\,.$ (4)
In addition to the intrinsic vacuum effect of the added mass eigenstate on the oscillation probabilities, matter effects are critically important in the 3+1 model. When active neutrinos propagate through a medium, they experience the weak potential induced by the nucleons and electrons within the matter. This leads to coherent forward scattering wolfensteinNeutrinoOscillationsMatter1978 , which can occur for all active flavours via neutral current (NC) interactions with neutrons, protons and electrons, and for $\nu_{e}$ or $\bar{\nu}_{e}$ via charged current (CC) interactions with electrons. For the CC channel, the corresponding effective potential is giuntiFundamentalsNeutrinoPhysics2007 :
$V_{\textrm{CC}}=\pm\sqrt{2}G_{F}N_{e}\,$, (5)
where $G_{F}$ is the Fermi constant, $N_{e}$ the electron density of the medium and $V_{\textrm{CC}}$ is positive for $\nu_{e}$, negative for $\bar{\nu}_{e}$. Neutral current interactions with neutrons, protons and electrons are equally possible for $\nu_{e}$, $\nu_{\mu}$ and $\nu_{\tau}$. The corresponding potential is:
$V_{\textrm{NC}}=V_{\textrm{NC}}^{n}+V_{\textrm{NC}}^{p}+V_{\textrm{NC}}^{e^{-}}$ (6)
where, due to equal densities of protons and electrons in ordinary matter and to their opposite charge, the potentials $V_{\textrm{NC}}^{p}$ and $V_{\textrm{NC}}^{e^{-}}$ cancel out. Only the neutron-induced potential remains, yielding:
$V_{\textrm{NC}}=\mp\frac{1}{2}\sqrt{2}G_{F}N_{n}$ (7)
where $N_{n}$ is the neutron density of the medium and $V_{\textrm{NC}}$ is negative for $\nu$, positive for $\bar{\nu}$.
In the standard three-flavour case, matter effects arise as only electron (anti)neutrinos are affected by $V_{\textrm{CC}}$, impacting their propagation and altering the oscillation patterns with respect to the vacuum case. However, all three flavours experience the same NC potential from neutrons, so no flavour-dependent coherent forward scattering is introduced by $V_{\textrm{NC}}$. Conversely, in the 3++1 case, sterile neutrinos are not affected by $V_{\textrm{NC}}$, with the neutrino propagation being described by the following Hamiltonian in the flavour basis:
$H_{\textrm{fl}}=UH_{0}U^{\dagger}+V_{4\nu}$ (8)
where $U$ is the 3++1 PMNS matrix in vacuum, $H_{0}=\frac{1}{2E}\textrm{diag}(0,\Delta m^{2}_{21},$ $\Delta m^{2}_{31},\Delta m^{2}_{41})$ is the Hamiltonian in vacuum in the mass basis and $V_{4\nu}=\textrm{diag}(V_{\textrm{CC}},0,0,$ $V_{\textrm{NC}})$, where $V_{\textrm{NC}}$ has been subtracted from the diagonal.
[…]
Figure 1: Difference in muon neutrino (left) and antineutrino (right) survival probability between the standard three-flavour model and a 3++1 model with $\Delta m^{2}_{41}=1$ eV2 and $\sin^{2}\theta_{24}=0.1$ for up-going neutrinos traveling through the Earth. Normal ordering is assumed and the standard oscillation parameter values are fixed to the NuFit 5.2 ones estebanFateHintsUpdated2020 . The other sterile parameters are set to 0. The L/E lines are shown in units of km/GeV. The oscillograms presented in this article are all computed with OscProb joao_coelho_2023_10104847 using the Preliminary Reference Earth Model dziewonskiPreliminaryReferenceEarth1981 .
The TeV $\bar{\nu}_{\mu}$ disappearance resonance was established more than twenty years ago as a distinctive signature of an eV-scale sterile neutrino in atmospheric neutrino oscillations nunokawaProbingLSNDMass2003 . As seen in figure 1, the effect is more prominent at $E_{\nu}\simeq 5$ TeV for mantle-only crossing trajectories ($\cos\theta>-0.838$). It can be interpreted as the consequence of the Mikheyev-Smirnov-Wolfenstein effect wolfensteinNeutrinoOscillationsMatter1978 ; ermilovav.k.BuildupNeutrinoOscillations1986 ; mikheyevResonantNeutrinoOscillations1989 in the Earth’s mantle. The $\bar{\nu}_{\mu}$ disappearance peak corresponds to a conversion into an almost pure sterile state bailly-salinsAtmosphericMuonStudies2024 . This feature has been extensively used in analyses of IceCube data to constrain the $\theta_{24}$ mixing angle as a function of $\Delta m_{41}^{2}$ razzaqueSearchingSterileNeutrinos2011 ; razzaqueSearchesSterileNeutrinos2012 ; esmailiConstrainingSterileNeutrinos2012 ; aartsenSearchingEVscaleSterile2020 ; icecubecollaborationSearchEVscaleSterile2024a . For KM3NeT/ORCA, the statistics of TeV-scale neutrinos are limited because of its comparatively smaller detector size.
Instead, the analysis presented in this paper primarily focuses on the lower energy feature highlighted in figure 1 (left). While the average effect of the presence of the sterile neutrino is an overall decrease of the $\nu_{\mu}$ survival probability compared to the three-flavour case, there is a notable exception. For the most vertical neutrinos, crossing both the mantle and the core ($\cos\theta<-0.838$), there is a region around $25$ GeV where less muon neutrinos disappear in the 3++1 case at the first $\nu_{\mu}\rightarrow\nu_{\tau}$ standard oscillation maximum. The effect of the $\theta_{24}$ and $\theta_{34}$ mixing angles (assuming $\delta_{24}=0$) on that feature is illustrated in the left panel of figure 2. As previously observed, when $\theta_{24}\neq 0$, the amplitude of the survival minimum is reduced. $\theta_{34}\neq 0$ has the same effect. The degeneracy between $\theta_{24}$ and $\theta_{34}$ can be partially lifted by the lower (average) $\nu_{\mu}$ survival probability outside the minimum region seen in the case $\theta_{24}\neq 0$ but not for $\theta_{34}\neq 0$. In addition, this figure shows that when both active-sterile mixing angles are non-zero, the position of the minimum can be significantly shifted. This suggests that KM3NeT/ORCA could be particularly sensitive to simultaneously large values of $\theta_{24}$ and $\theta_{34}$. The right panel of figure 2 illustrates that the $\delta_{24}$ phase also affects the position of the minimum, with a limited effect on its amplitude (compared to the mixing angles). The known degeneracy between the NMO and the sign of $\cos\delta_{24}$ aielloSensitivityLightSterile2021 is also illustrated in figure 2 (right) as the curve obtained assuming IO and $\delta_{24}=0$ overlaps the $\delta_{24}=\pi$ curve obtained in NO over most of the energy range considered. As all the effects mentioned here affect the $\nu_{\mu}\rightarrow\nu_{\tau}$ oscillations, degeneracies with the standard atmospheric parameters $\Delta m^{2}_{31}$ and $\theta_{23}$ are also expected. For instance, an apparent shift in $\Delta m^{2}_{31}$ could be accommodated by non-zero values for $\theta_{24}$ and $\theta_{34}$.
Figure 2: Muon neutrino survival probability as a function of $E_{\nu}$, for vertically up-going neutrinos ($\cos\theta=-1.0$). Normal ordering is assumed except when mentioned otherwise, with NuFit 5.2 values estebanFateHintsUpdated2020 for the standard oscillation parameters. The three-flavour model (dashed black curve) is compared with various 3++1 models all assuming $\Delta m^{2}_{41}=1$ eV2. Left: different values of $\theta_{24}$ and $\theta_{34}$ ($s^{2}_{ij}=\sin^{2}\theta_{ij}$), with $\delta_{24}$ and other active-sterile mixing parameters fixed to 0. Right: different $\delta_{24}$ values, with $\sin^{2}\theta_{24}=0.07$, $\sin^{2}\theta_{34}=0.1$ and other active-sterile mixing parameters fixed to 0.
3 The KM3NeT/ORCA detector
The KM3NeT collaboration is building two water Cherenkov neutrino telescopes at the bottom of the Mediterranean Sea adrian-martinezLetterIntentKM3NeT2016 . The KM3NeT/ARCA (Astroparticle Research with Cosmics in the Abyss) detector, optimised for the detection of neutrinos from astrophysical sources in the TeV to PeV energy range, is located at a depth of about 3500 m, 80 km offshore Portopalo di Capo Passero in Sicily. The KM3NeT/ORCA (Oscillation Research with Cosmics in the Abyss) detector, optimised for the determination of the NMO through the measurement of oscillations of GeV to TeV atmospheric neutrinos crossing the Earth, is located 40 km offshore Toulon (France) at a depth of about 2500 m. Oscillation studies in the KM3NeT/ORCA energy range are well-suited to probe Beyond the Standard Model (BSM) hypotheses such as the 3++1 sterile neutrino model.
Both detectors collect the Cherenkov light induced by the relativistic charged particles emerging from neutrino interactions. They consist of 3-dimensional arrays of glass spheres named Digital Optical Modules (DOMs) aielloKM3NeTMultiPMTOptical2022a housing 31 3-inch photomultiplier tubes (PMTs) each, arranged along flexible vertical structures called Detection Units (DUs) carrying 18 DOMs each. For KM3NeT/ORCA, the vertical spacing between DOMs is $9$ m. Each DU is anchored to the seabed and remains vertical due to the buoyancy of the DOMs and of a buoy installed at its top. The DUs are arranged following a cylindrical footprint, with a horizontal spacing of $20$ m. When completed, KM3NeT/ORCA will consist of 115 DUs, instrumenting a mass of sea water of around $7$ Mton.
When the pulse resulting from the detection of a Cherenkov photon by a PMT reaches a certain threshold, a hit is generated. To save bandwidth, rather than digitizing the whole pulse, only the Time Over Threshold (ToT) of the pulse is recorded, along with the time of the hit and the PMT identifier. The hit information is sent to the onshore data processing center for online data filtering. More information about the data acquisition system of KM3NeT can be found in chiarusiKM3NeTDataAcquisition2023 . Events are generated by trigger algorithms looking for causally-related hits on multiple DOMs. The trigger algorithms are designed to suppress background hits coming from dark counts (spontaneous electron emission from the photocathode of the PMTs) aielloCharacterisationHamamatsuPhotomultipliers2018 , bioluminescence haddockBioluminescenceSea2010 ; widderBioluminescencePelagicVisual2002a ; aguzziInertialBioluminescenceRhythms2017 , and natural radioactivity of 40K present in the sea water km3netcollaborationDependenceAtmosphericMuon2020 . The KM3NeT detectors are not sensitive to the detailed particle content of an event. Events are grouped into two topologies in KM3NeT/ORCA: track-like and shower-like events. Track-like events are caused by muons crossing the detector, encompassing $\nu_{\mu}$-CC interactions and $\nu_{\tau}$-CC interactions quickly followed by the decay of the $\tau$ lepton into a muon (branching ratio $\simeq 17.4\%$ particledatagroupReviewParticlePhysics2024 ). Track-like events also include large amounts of down-going high-energy atmospheric muons which still reach the detectors despite the natural shielding provided by the sea water aielloAtmosphericMuonsMeasured2024 . GeV muons travel in a straight line and lose energy continuously at a minimum rate of about 0.2 GeV/m. Hence, the induced hits are observed on DOMs distributed evenly inside a cylinder, whose axis is the muon trajectory. The other neutrino interaction channels ($\nu_{e}$-$C$, other $\nu_{\tau}$-$C$, $\nu$-NC) create particle cascades (hadronic and/or electromagnetic), and are hence shower-like events. They appear as a localised burst of light in the detector, with hits distributed inside an ellipsoid whose longitudinal extent is 10 m at most adrian-martinezLetterIntentKM3NeT2016 . Based on the expected hit distributions, three trigger algorithms are applied in parallel: one looking for track-like events as sets of causally related hits within the volume of a cylinder, and two looking for shower-like events as sets of causally related hits within the volume of a sphere, with one of the two optimised for low-energy events.
Events are reconstructed assuming either a track or a shower model. More information about the track and shower reconstruction is given in refs. ofearraighFollowingLightNovel2024 and domiShowerReconstructionSterile2019a respectively. Both procedures are iterative and use a maximum likelihood estimation based on the multi-dimensional probability density function of the photon arrival time on the PMTs.
To fulfill the physics goals of KM3NeT, good angular and energy resolutions must be achieved. Both require an accurate calibration of the detector. The KM3NeT detectors are huge infrastructures deployed in a changing environment which affects their performance over time. The position and orientation of the DOMs are measured every 10 minutes with an accuracy of 20 cm using an acoustic positioning system and sets of accelerometers and magnetometers located inside each DOM gatiusoliverDynamicalPositionOrientation2023 . The synchronisation of all PMTs with an onshore master clock is ensured through a custom implementation of the White Rabbit protocol serranoWhiteRabbitProject2009 ; aielloKM3NeTBroadcastOptical2023 . As delays can occur at multiple levels of the infrastructure, the time calibration must be achieved at various scales: between the PMTs of a given DOM (intra-DOM), between the DOMs of a given DU (inter-DOM), and between the DUs (inter-DU). The intra-DOM calibration uses coincident hits induced by 40K decays melisInSituCalibrationKM3NeT2017 ; the inter-DOM calibration uses dedicated LED flashers called nanobeacons aielloNanobeaconTimeCalibration2022 ; and the inter-DU calibration relies on the maximization of the quality of the reconstructed tracks of atmospheric muons bailly-salinsAtmosphericMuonStudies2024 ; bailly-salinsTimePositionOrientation2023 .
4 ORCA6 data sample and event selection
Data were taken with the ORCA6 configuration from January 2020 to November 2021, with a total detector livetime of 633 days. After applying the data quality cuts presented in aielloMeasurementNeutrinoOscillation2024 , a total of 510 days of data taking remains, corresponding to a total exposure of 433 kton-years. This data sample was also used for other recently published ORCA6 oscillation analyses, comprising a standard neutrino oscillation analysis aielloMeasurementNeutrinoOscillation2024 , a study of tau neutrinos and unitarity aielloStudyTauNeutrinos2025 , searches for non-standard neutrino interactions aielloSearchNonstandardNeutrino2025a , invisible neutrino decay aielloProbingInvisibleNeutrino2025a and quantum decoherence aielloSearchQuantumDecoherence2025 .
In order to remove accidental triggers, pre-selection cuts on the number of hits and the reconstructed track likelihood of the events are applied aielloMeasurementNeutrinoOscillation2024 . In addition, only events reconstructed as up-going are kept. Even though only a small fraction of atmospheric muons are misreconstructed as up-going, they constitute close to $99.9\%$ of the accepted events at this stage, as the rate of reconstructed atmospheric muon events ($\approx 7$ Hz) exceeds that of neutrino events by 5 orders of magnitude.
The event selection and classification is performed using two dedicated sets of Boosted Decision Trees (BDTs), trained using Monte Carlo (MC) simulations (see section 5.1). Both BDTs are trained on (and applied to) events passing the pre-selection cuts only. The BDT models are built using variables related to the space and time distribution of the hits, as well as variables computed through the track and shower reconstruction chains, which are both applied to all triggered events. Each BDT model summarises the input features into a single output score. Details on the training and performances of the BDTs are given in aielloMeasurementNeutrinoOscillation2024 . The first BDT is used to separate neutrino events from the misreconstructed atmospheric muons events. It is able to identify poorly reconstructed atmospheric muons because they trigger more hits in the upper hemisphere of the DOMs and in the upper part of the DUs. Events with a high atmospheric muon BDT score are removed, resulting in a neutrino sample that contains 5828 events, with an expected muon contamination of a few percent. The goal of the second BDT is the separation between track-like and shower-like neutrino events. The resulting track sample is further divided in two classes: a high purity track class with an expected atmospheric muon contamination around $0.1\%$, and around $95\%$ $\nu_{\mu}$-CC events; and a low purity track class with a few percent muon contamination and around $90\%$ $\nu_{\mu}$-CC events. This selection leads to a total of 1868, 2002 and 1958 events in the high purity track, low purity track and shower classes respectively.
The track/shower separation gives some sensitivity to the flavour of the interacting neutrino, as events classified as tracks are mostly $\nu_{\mu}$-CC. However, around $45\%$ of the selected shower-like events are low-energy (<10<10 GeV) $\nu_{\mu}$-CC events, as short muon tracks cannot be easily distinguished from shower-like events. The low/high purity separation for track-like events increases the sensitivity to the standard oscillation parameters by isolating tracks with the best angular resolution in the high purity class aielloMeasurementNeutrinoOscillation2024 . For the eV-scale sterile neutrino search presented here, the same class definitions are kept, as the main feature giving sensitivity to the active-sterile mixing angles of interest is the first $\nu_{\mu}\rightarrow\nu_{\tau}$ oscillation maximum (see section 2), the same used to measure $\theta_{23}$ and $\Delta m^{2}_{31}$ aielloMeasurementNeutrinoOscillation2024 . The reconstructed energy range for the two track classes is 2 GeV to 100 GeV. It is extended to 1 TeV for the shower class.
5 Analysis method
This analysis relies on comparing the observed numbers of reconstructed and selected events to the expected numbers, given a set of oscillation parameters of interest and nuisance parameters. This comparison is performed as a function of the reconstructed neutrino energy $E$ and the cosine of the zenith angle $\cos\theta$.
5.1 Event distribution modeling and Monte Carlo simulations
To compute the expected number of events, the atmospheric neutrino flux is multiplied by the oscillation probabilities and weighted by the CC and NC interaction cross-sections, before accounting for the detector response. For the atmospheric neutrino flux, the year-averaged, solar minimum HKKM 2014 hondaAtmosphericNeutrinoFlux2015 flux tables computed for the Fréjus experiment site (located in Modane, 250 km away from the KM3NeT/ORCA site) are used. The OscProb software joao_coelho_2023_10104847 is used to compute the oscillation probabilities, and the Earth density is described using the Preliminary Reference Earth Model dziewonskiPreliminaryReferenceEarth1981 . The detector response is modeled through response matrices computed from MC simulations, briefly described in the following (see adrian-martinezLetterIntentKM3NeT2016 and aielloDeterminingNeutrinoMass2022 for more details). The generation of neutrino events is performed using the GENIE-based andreopoulosGENIENeutrinoMonte2010 software gSeaGen aielloGSeaGenKM3NeTGENIEbased2020 . Atmospheric muons are generated using the parametric generator MUPAGE carminatiAtmosphericMUonsPArametric2008 . The propagation of the Cherenkov photons induced along the path of the produced charged particles is performed using the Geant4-based agostinelliGeant4aSimulationToolkit2003 package KM3Sim tsirigotisHOUReconstructionSimulation2011 . For high-energy particles, a faster, custom KM3NeT package relying on probability density functions of the light arrival time is used instead of KM3Sim because of the large amounts of photons generated. The PMT response is simulated using another custom KM3NeT package which also includes the simulation of dark count and optical background, based on the trigger rates observed in the data for each PMT. From this point on, the output of the MC simulation chain is processed using the same trigger and event reconstruction algorithms (see section 3) and the same selection and classification procedure (see section 4) as for data. A 4-dimensional response matrix is built for each event class and each neutrino interaction channel, mapping the true neutrino energy and cosine of the zenith angle to the reconstructed ones. The choice of the binning used to build the response matrices is described in aielloMeasurementNeutrinoOscillation2024 . The full chain to compute response matrices and expected event distributions is implemented through the custom KM3NeT software Swim bourretNeutrinoOscillationsEarth2018 ; carreterocuencaNeutrinoOscillationsInvisible2024 .
5.2 Nuisance parameters
To model sources of systematic uncertainties, a total of 16 nuisance parameters are used in this analysis. Most are common to the standard oscillation analysis aielloMeasurementNeutrinoOscillation2024 :
  • Uncertainties on the atmospheric neutrino flux shape and neutrino species ratios are modeled through corrections applied directly to the HKKM flux: on the ratio of up-going to horizontal neutrino ($\delta_{\theta}$), on the spectral index ($\delta_{\sigma}$), and on the ratios of, respectively, $\nu_{\mu}$ to $\bar{\nu}_{\mu}$ ($s_{\mu\bar{\mu}}$), $\nu_{e}$ to $\bar{\nu}_{e}$ ($s_{e\bar{e}}$), and $\nu_{e}+\bar{\nu}_{e}$ to $\nu_{\mu}+\bar{\nu}_{\mu}$ ($s_{e\mu}$).
  • Uncertainties on neutrino cross-sections and on the selection and classification efficiencies are modeled through scaling factors affecting the normalisation of specific channels. The overall normalisation $f_{\textrm{all}}$ scales all selected events. The normalisation of the NC ($\tau$-$C$) events $f_{\textrm{NC}}$ ($f_{\tau\textrm{CC}}$) accounts for the uncertainty in modelling the NC interaction ($\tau$-$C$) cross-section and event selection. The $f_{\textrm{HPT}}$ and $f_{\textrm{S}}$ normalisations are used for the relative normalisation of the high purity track and shower classes. A normalisation for the atmospheric muon background $f_{\mu}$ is also used. An additional scaling factor $f_{\textrm{HE}}$ is used to scale high-energy events to account for the different assumptions made on light propagation by the two different light propagation software packages mentioned in section 5.1. This scaling is applied for NC events with true energy above 100 GeV and for CC events with true energy above 500 GeV.
  • Uncertainties on the water absorption and scattering lengths and the absolute PMT efficiency are accounted for through an energy scale parameter $E_{s}$. This parameter is applied as a shift in the true energy of the detector response. For more details and discussion about the implementation, see chauStudyAtmosphericNeutrino2021 .
For the sterile neutrino search, $\Delta m^{2}_{31}$ and $\theta_{23}$ are also nuisance parameters. No prior uncertainty is assumed for $\theta_{23}$. $\Delta m^{2}_{31}$ is constrained using the values and uncertainties reported by the Daya Bay collaboration dayabaycollaborationPrecisionMeasurementReactor2023 . The Daya Bay measurement, obtained from $\nu_{e}$ disappearance, can be considered uncorrelated with the sterile neutrino search presented in this work, which relies mostly on the $\nu_{\mu}$ disappearance channel in a “no $\nu_{e}$” approximation (see section 6). The other standard oscillation parameters, for which KM3NeT/ORCA has no sensitivity, are fixed to the NuFit 5.0 values with Super-Kamiokande data included in the global fit estebanFateHintsUpdated2020 given in table 1. Finally, the $\delta_{24}$ phase which, as shown in figure 2, influences the position of the $\nu_{\mu}\rightarrow\nu_{\tau}$ standard oscillation maximum, is fitted without prior constraint.
Parameter NO IO
$\theta_{12}$ $^{\circ} 33.4433.44^{\circ}$
$\theta_{13}$ $8.57^{\circ}$ $8.60^{\circ}$
$\Delta m_{21}^{2}$ $10^{-5}~\mathrm{eV^{2}}$] $7.42$ 7.42
$\delta_{\text{CP}}$ $197^{\circ}$ $282^{\circ}$
Table 1: Values used for the non-fitted standard neutrino oscillation parameters, from NuFit 5.0 including Super-Kamiokande data estebanFateHintsUpdated2020 .
5.3 Statistical methods
In the following, given an observed number of events $n_{ij}$ of the class $i$ in bin $j$, the parameters of interest of the model $\vec{x}$ are determined through the minimisation of the following negative log-likelihood function:
$-2\ln\mathcal{L}(\vec{x},\vec{\eta})=2$ $\sum_{i=1}^{N_{\textrm{classes}}}\sum_{j=1}^{N_{\textrm{bins}}}\left[\beta_{ij}\mu_{ij}(\vec{x},\vec{\eta})-n_{ij}+n_{ij}\ln\left(\frac{n_{ij}}{\beta_{ij}\mu_{ij}(\vec{x},\vec{\eta})}\right)+\frac{(\beta_{ij}-1)^{2}}{\sigma_{\beta_{ij}}^{2}}\right]$ (9)
$+$ $\sum_{k=1}\left(\frac{\eta_{k}-\langle\eta_{k}\rangle}{\sigma_{k}}\right)^{2}\,$,
where $\mu_{ij}(\vec{x},\vec{\eta})$ is the expected event distribution for the set of parameters of interest $\vec{x}$ and nuisance parameters $\vec{\eta}$. The $\beta_{ij}$ factors act as additional nuisance parameters and are introduced to account for the finite number of generated MC events used to build the response matrices, following the Barlow and Beeston light method barlowFittingUsingFinite1993 ; conwayIncorporatingNuisanceParameters2011 (see ref carreterocuencaNeutrinoOscillationsInvisible2024 for details). The term in square brackets corresponds to a Poissonian likelihood. The term in parentheses is a Gaussian penalisation term accounting for the prior uncertainties on the subset of constrained nuisance parameters: for a parameter $\eta_{k}$, $\langle\eta_{k}\rangle$ is the prior mean and $\sigma_{k}$ is its standard deviation. In Swim, the parameters of interest $\vec{x}$ and the nuisance parameters $\vec{\eta}$ are determined by minimising equation 9 using the MIGRAD solver of the Minuit2 package ROOTMinuit2MnMigrad .
The test statistic used to build the confidence interval of the parameters of interest around their best-fit values is the logarithm of the likelihood ratio between each point of the phase-space of interest $\vec{x}$ and the best-fit point $\vec{\widehat{x}}$, written $2\Delta\ln\mathcal{L}$ in the following. The confidence intervals of the parameters of interest are built in a frequentist way. Using Wilks’ theorem wilksLargeSampleDistributionLikelihood1938 , the allowed region of the parameter space given a certain confidence level is obtained by comparing the value of $2\Delta\ln\mathcal{L}$ with the corresponding quantile of a $\chi^{2}$ distribution with two degrees of freedom, as two is the dimension of the parameter space of interest.
5.4 Fitting procedure
For each point of the phase space of interest, equation 9 is minimised using 2 starting points for 4 nuisance parameters, to avoid getting stuck in local minima: normal and inverted ordering ($\Delta m^{2}_{31}=\{2.541\times 10^{-3},-2.496\times 10^{-3}\}~\mathrm{eV^{2}}$); $\theta_{23}$ lower and upper octant ($\theta_{23}=\{40^{\circ},50^{\circ}\}$); energy scale $E_{s}$ below and above 1.0 ($E_{s}=\{0.95,1.05\}$); and $\delta_{24}$ below and above $180^{\circ}$ ($\delta_{24}=\{90^{\circ},270^{\circ}\}$). This produces a total of 16 distinct starting points. The two start values for $\Delta m_{31}^{2}$, $\theta_{23}$ and $E_{s}$ are common to the standard oscillation analysis of ORCA6 aielloMeasurementNeutrinoOscillation2024 . For $\delta_{24}$, the use of two start values is motivated by the observation of local minima when computing the $2\Delta\ln\mathcal{L}$ profile of that parameter. In each fit, the parameter space is restricted to the NMO, $\theta_{23}$ octant and $(E_{s}-1)$ sign corresponding to the start value. No limit is put on $\delta_{24}$.
6 Results
The ORCA6 data sample is analysed under the $\Delta m_{41}^{2}=1$ eV2 hypothesis, although the results also hold for $\Delta m_{41}^{2}>1$ eV2 for which faster oscillations are unobservable. Simultaneous constraints on the mixing of the $\nu_{\mu}$ and $\nu_{\tau}$ states with the sterile state are obtained by scanning the $(|U_{\mu 4}|^{2},|U_{\tau 4}|^{2})$ space, with $U_{e4}=0$ (i.e. $\theta_{14}=\delta_{14}=0$). In that case, the magnitude of the $U_{\mu 4}$ and $U_{\tau 4}$ mixing elements is expressed as:
$|U_{\mu 4}|^{2}$ $=\,\sin^{2}\theta_{24}$ (10)
$|U_{\tau 4}|^{2}$ $=\,\cos^{2}\theta_{24}\sin^{2}\theta_{34}\,.$ (11)
For each fit $\delta_{24}$ is set free. To obtain the allowed region for the parameters of interest at a given confidence level, a 2-dimensional scan is performed by sampling the $(|U_{\mu 4}|^{2},|U_{\tau 4}|^{2})$ phase space uniformly in log scale, using 27 points between $1\times 10^{-3}$ and $0.5$ (excluded) for both $|U_{\mu 4}|^{2}$ and $|U_{\tau 4}|^{2}$, with an additional point at 0.
6.1 Best-fit point
The position of the best-fit point $\vec{\widehat{x}}$ is computed independently of the grid scan: fits with free $U_{\mu 4}$ and $U_{\tau 4}$ (thus free $\theta_{24}$ and $\theta_{34}$) are performed. To avoid falling in potential local minima in the $(|U_{\mu 4}|^{2},|U_{\tau 4}|^{2})$ space, three pairs of starting values are used (in addition to the $16$ sets of starting values already introduced): $(0.0,0.0)$, $(0.5,0.0)$ and $(0.0,0.5)$. This makes a total of 48 fits to determine the best-fit point, found at:
[…]
Parameter Nominal value $\pm$ uncertainty Best Fit Post-fit uncertainty
$\delta_{\theta}$ 0.00 $\pm$ 0.02 -0.00 -0.02/+0.02
$\delta_{\gamma}$ 0.00 $\pm$ 0.30 -0.00 -0.03/+0.03
$s_{e\bar{e}}$ 0.00 $\pm$ 0.07 0.01 -0.07/+0.07
$s_{\mu\bar{\mu}}$ 0.00 $\pm$ 0.05 0.00 -0.05/+0.05
$s_{e\mu}$ 0.00 $\pm$ 0.02 -0.00 -0.02/+0.02
$f_{\tau\textrm{CC}}$ 1.00 $\pm$ 0.20 0.90 -0.18/+0.18
$f_{\textrm{NC}}$ 1.00 $\pm$ 0.20 0.85 -0.19/+0.19
$E_{s}$ 1.00 $\pm$ 0.09 1.10 -0.07/+0.07
$f_{\textrm{HE}}$ $1.00\pm 0.50$ 1.57 -0.29/+0.32
$f_{\textrm{HPT}}$ 1.00 0.92 -0.04/+0.04
$f_{\textrm{S}}$ 1.00 0.89 -0.06/+0.06
$f_{\mu}$ 1.00 0.40 -0.31/+0.36
$f_{\textrm{all}}$ 1.00 1.27 -0.14/+0.15
$\Delta m_{31}^{2}$ $10^{-3}$ eV2] $2.541\pm 0.060$ 2.521 -0.059/+0.059
$\theta_{23}$ [°] 49.2 44.1 -4.1/+6.4
$\delta_{24}$ [°] 0 322 -322/+38
Table 2: Best-fit values and post-fit uncertainties at $68\%$ CL of the nuisance parameters from the fit of ORCA6 data to $U_{\mu 4}$ and $U_{\tau 4}$ with $\Delta m_{41}^{2}=1$ eV2. The fit is done in both mass orderings, and the best fit is found in NO.
6.2 $(|U_{\mu 4}|^{2},|U_{\tau 4}|^{2})$ scan results
The log-likelihood ratio map resulting from the 2-dimensional grid scan, as well as upper limits of the allowed parameter space at confidence levels (CL) of $90\%$, $95\%$ and $99\%$, are shown in figure 3 (left). A linear scale is used for the $y$ axis so that the best fit, which is outside the scanned range, is visible. The $(0,0)$ point, with a value $2\Delta\ln\mathcal{L}(0,0)=1.76$, is within the allowed region for the three CLs considered. This means that the ORCA6 data is fully compatible with the standard model. When profiling over the other mixing element, the upper limits at a $90\%$ CL are found to be:
$|U_{\mu 4}|^{2}$ < $<0138$ (13)
$|U_{\tau 4}|^{2}$ < $<0076\,.$
Figure 3: Left: Log-likelihood ratio map obtained from the ORCA6 data sample over the $|U_{\mu 4}|^{2}$ and $|U_{\tau 4}|^{2}$ phase space. $2\Delta\log\mathcal{L}$ is not displayed when higher than $10$ units. The black lines show upper limits of the allowed parameter space at various confidence levels. Right: Comparison of the observed and expected upper limits at $90\%$ CL on the $U_{\mu 4}$ and $U_{\tau 4}$ mixing elements for ORCA6. The excluded regions are on the top right side of the contours.
In figure 3 (right), the $90\%$ CL upper limits in $|U_{\mu 4}|^{2}$ and $|U_{\tau 4}|^{2}$ obtained by fitting the ORCA6 433 kton-years data set are compared with the ones obtained by fitting an Asimov MC data set generated using the data best-fit point values for the parameters of interest and the nuisance parameters (reported in table 2). This comparison shows that a larger region of the phase space is excluded when fitting the data, compared to the one expected from the Asimov sensitivity study. The stricter upper limits obtained when fitting the data can be understood by looking at the one-dimensional reconstructed $L/E$ event distributions of figure 4, shown for the three event classes. Both the data events and model best-fit point for this analysis are shown. The standard oscillation best fit is also shown for reference. It is barely distinguishable from the sterile best fit. The standard oscillation pattern is clearly visible in the high purity (and to a lower extent, low purity) track class. However, it is hardly distinguishable in the shower class, due to the mixture of flavours in this class. On the high purity tracks distribution, the data underfluctuates with respect to the model at the oscillation dip (corresponding to the first $\nu_{\mu}\rightarrow\nu_{\tau}$ maximum) at $L/E\sim\mathcal{O}(10^{3})$ km/GeV. This is a crucial feature of the event distribution studied here as it constrains the model to give the deepest dip possible, which strongly favours maximal mixing for $\theta_{23}$ (as already reported in aielloMeasurementNeutrinoOscillation2024 ). It also puts strong constraints on the active-sterile mixing angles $\theta_{24}$ and $\theta_{34}$, with particularly strict limits on $\theta_{34}$. Indeed, the effect of $\theta_{34}$ on the $P(\nu_{\mu}\rightarrow\nu_{\mu})$ first maximum is very similar to the effect of a $\theta_{23}$ value far from maximal mixing, as illustrated in figure 5, while $\theta_{24}$ lowers the survival probability outside the oscillation dip region. This explains why high values of $\theta_{34}$ are strongly rejected. For illustration, an arbitrarily high value of $|U_{\tau 4}|^{2}$ is chosen to draw the dashed orange curve in figure 4, showing that higher $\theta_{34}$ makes the oscillation dip shallower, and therefore even further away from the data.
Figure 4: Event distributions in $L/E$ for each class of the ORCA6 data (black), compared with the model prediction at the best-fit point in the eV sterile neutrino case (red) and in the standard oscillation case (dashed blue). The dashed orange curve shows the model prediction for an arbitrary point in the $(|U_{\mu 4}|^{2},|U_{\tau 4}|^{2})$ phase space.
Figure 5: Muon neutrino survival probability as a function of the neutrino energy under various scenarios for $\cos\theta=-1.0$. The three-flavour standard oscillation model assuming the nominal value for $\theta_{23}=49.1\degree$ (dashed black) is compared with the three-flavour model with $\theta_{23}=59\degree$ (dashed grey) and with the 3++1 models with $\sin^{2}\theta_{24}=0.1$ (pink) and $\sin^{2}\theta_{34}=0.15$ (yellow), for $\Delta m_{41}^{2}=1$ eV2. All other sterile parameters are set to 0.
6.3 Comparison with existing measurements
The ORCA6 upper limits at $90\%$ CL are compared with the results of other experiments in figure 6. They are competitive with these existing results, especially concerning the limit on the $\nu_{\tau}$ mixing with the sterile state: only DeepCore, using $7.5$ years of data, puts a stronger constraint on $|U_{\tau 4}|^{2}$. For $|U_{\mu 4}|^{2}$, ANTARES, IceCube, DeepCore and Super-Kamiokande all put stronger constraints than ORCA6, but these results were obtained with several years of data-taking from complete detectors, while the ORCA6 measurement was obtained with 1.4 years of data recorded with a detector that was $5\%$ of its final size. All analyses considered here treat $\delta_{24}$ as a free nuisance parameter, with the exception of Super-Kamiokande and IceCube. The Super-Kamiokande analysis fixes $\delta_{24}=0$, leading to more restrictive limits than with a free $\delta_{24}$. The IceCube analysis relies on the TeV disappearance feature which is less sensitive to $\delta_{24}$. It uses $\delta_{24}=\pi$, as this choice was found to yield the most conservative limits. The performance of KM3NeT/ORCA is due to its good sensitivity to $\nu_{\mu}$ in the 203020-30 GeV region which allows for the monitoring of the first $\nu_{\mu}\rightarrow\nu_{\tau}$ standard oscillation maximum. Other detectors have to rely on weaker effects at lower energy (e.g. Super-Kamiokande) or are optimised for slightly higher energies (e.g. ANTARES).
Figure 6: Comparison of the upper limits at $90\%$ CL on $|U_{\mu 4}|^{2}$ and $|U_{\tau 4}|^{2}$ obtained by ORCA6 with previous measurements from Super-Kamiokande super-kamiokandecollaborationLimitsSterileNeutrino2015 , NOvA thenovacollaborationSearchActiveSterileAntineutrino2021 , ANTARES albertMeasuringAtmosphericNeutrino2019 , IceCube abbasiExplorationMassSplitting2024a and DeepCore icecubecollaborationSearchLightSterile2024 . The excluded regions are on the top right side of the contours.
7 Conclusion
The first search for an eV-scale sterile neutrino performed with the KM3NeT/ORCA detector is presented in this article. In this analysis, a data sample collected with 6 detection units ($5\%$ of the complete detector size) is used, corresponding to an exposure of 433 kton-years and containing 5828 neutrino candidates. Constraints on the mixing of a hypothetical eV-scale sterile neutrino with the $\nu_{\mu}$ and $\nu_{\tau}$ states are established under the assumption $\Delta m^{2}_{41}=1$ eV2. The results are compatible with the three-flavour standard model, and the upper limits obtained on $|U_{\mu 4}|^{2}$ and $|U_{\tau 4}|^{2}$ are compatible with the results reported by other experiments.
This result, although based on a limited detector size and statistics, highlights the significant potential of KM3NeT/ORCA. More specifically for eV-scale sterile neutrino searches, it should be possible to probe the sterile-induced TeV resonance on $\bar{\nu}_{\mu}$ disappearance with the final configuration of KM3NeT/ORCA.
Acknowledgements.
The authors acknowledge the financial support of: KM3NeT-INFRADEV2 project, funded by the European Union Horizon Europe Research and Innovation Programme under grant agreement No 101079679; Funds for Scientific Research (FRS-FNRS), Francqui foundation, BAEF foundation. Czech Science Foundation (GAČR 24-12702S); Agence Nationale de la Recherche (contract ANR-15-CE31-0020), Centre National de la Recherche Scientifique (CNRS), Commission Européenne (FEDER fund and Marie Curie Program), LabEx UnivEarthS (ANR-10-LABX-0023 and ANR-18-IDEX-0001), Paris Île-de-France Region, Normandy Region (Alpha, Blue-waves and Neptune), France, The Provence-Alpes-Côte d’Azur Delegation for Research and Innovation (DRARI), the Provence-Alpes-Côte d’Azur region, the Bouches-du-Rhône Departmental Council, the Metropolis of Aix-Marseille Provence and the City of Marseille through the CPER 2021-2027 NEUMED project, The CNRS Institut National de Physique Nucléaire et de Physique des Particules (IN2P3); Shota Rustaveli National Science Foundation of Georgia (SRNSFG, FR-22-13708), Georgia; This research was funded by the European Union (ERC MuSES project No 101142396); The General Secretariat of Research and Innovation (GSRI), Greece; Istituto Nazionale di Fisica Nucleare (INFN) and Ministero dell’Università e della Ricerca (MUR), through PRIN 2022 program (Grant PANTHEON 2022E2J4RK, Next Generation EU) and PON R&I program (Avviso n. 424 del 28 febbraio 2018, Progetto PACK-PIR01 00021), Italy; IDMAR project Po-Fesr Sicilian Region az. 1.5.1; A. De Benedittis, W. Idrissi Ibnsalih, M. Bendahman, A. Nayerhoda, G. Papalashvili, I. C. Rea, A. Simonelli have been supported by the Italian Ministero dell’Università e della Ricerca (MUR), Progetto CIR01 00021 (Avviso n. 2595 del 24 dicembre 2019); KM3NeT4RR MUR Project National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 3.1, Funded by the European Union – NextGenerationEU,CUP I57G21000040001, Concession Decree MUR No. n. Prot. 123 del 21/06/2022; Ministry of Higher Education, Scientific Research and Innovation, Morocco, and the Arab Fund for Economic and Social Development, Kuwait; Nederlandse organisatie voor Wetenschappelijk Onderzoek (NWO), the Netherlands; The grant ”AstroCeNT: Particle Astrophysics Science and Technology Centre”, carried out within the International Research Agendas programme of the Foundation for Polish Science financed by the European Union under the European Regional Development Fund; The program: ’Excellence initiative-research university’ for the AGH University in Krakow; The ARTIQ project: UMO-2021/01/2/ST6/00004 and ARTIQ/0004/2021; Ministry of Education and Scientific Research, Romania; Slovak Research and Development Agency under Contract No. APVV-22-0413; Ministry of Education, Research, Development and Youth of the Slovak Republic; MCIN for PID2021-124591NB-C41, -C42, -C43 and PDC2023-145913-I00 funded by MCIN/AEI/10.13039/501100011033 and by ”ERDF A way of making Europe”, for ASFAE/2022/014 and ASFAE/2022 /023 with funding from the EU NextGenerationEU (PRTR-C17.I01) and Generalitat Valenciana, for Grant AST22_6.2 with funding from Consejería de Universidad, Investigación e Innovación and Gobierno de España and European Union - NextGenerationEU, for CSIC-INFRA23013 and for CNS2023-144099, Generalitat Valenciana for CIDEGENT/2020/049, CIDEGENT/2021/23, CIDEIG/2023/20, ESGENT2024/24, CIPROM/2023/51, GRISOLIAP/2021/192 and INNVA1/2024/110 (IVACE+i), Spain; Khalifa University internal grants (ESIG-2023-008, RIG-2023-070 and RIG-2024-047), United Arab Emirates; The European Union’s Horizon 2020 Research and Innovation Programme (ChETEC-INFRA - Project no. 101008324). Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.